# Carré du champ flow matching: better quality-generalisation tradeoff in generative models

### v1

@@ Abstract
Deep generative models often face a fundamental tradeoff: high sample quality can come at the cost of memorisation, where the model reproduces training data rather than generalising across the underlying data geometry. We introduce Carré du champ flow matching (CDC-FM), a generalisation of flow matching (FM), that improves the quality-generalisation tradeoff by regularising the probability path with a geometry-aware noise. Our method replaces the homogeneous, isotropic noise in FM with a spatially varying, anisotropic Gaussian noise whose covariance captures the local geometry of the latent data manifold. We prove that this geometric noise can be optimally estimated from the data and is scalable to large data. Further, we provide an extensive experimental evaluation on diverse datasets (synthetic manifolds, point clouds, single-cell genomics, animal motion capture, and images) as well as various neural network architectures (MLPs, CNNs, and transformers). We demonstrate that CDC-FM consistently offers a better quality-generalisation tradeoff. We observe significant improvements over standard FM in data-scarce regimes and in highly non-uniformly sampled datasets, which are often encountered in AI for science applications. Our work provides a mathematical framework for studying the interplay between data geometry, generalisation and memorisation in generative models, as well as a robust and scalable algorithm that can be readily integrated into existing flow matching pipelines.
@@ 1 Introduction
Deep generative models aim to sample from an unknown probability density $\nu(x)$, given only finitely many training points $\mathcal{D}=\{x^{(i)}\}^{N}_{i=1}\subset\mathbb{R}^{d}$. Prominent paradigms include variational autoencoders (Kingma & Welling, 2014), generative adversarial networks (Mahdizadehaghdam et al., 2019), diffusion processes (Sohl-Dickstein et al., 2015; Ho et al., 2020), and methods based on continuous normalising flows (Chen et al., 2019; Song & Ermon, 2019; Lipman et al., 2023; Albergo et al., 2023) (CNFs). Among these, CNFs have had striking recent success across domains from image generation to molecule design and weather prediction, owing to their ability to generate high-quality samples. At the same time, quality alone cannot measure the goodness of the generative model, as high quality can be achieved by reproducing training points or close variants, known as memorisation, which has become a recurring concern (Somepalli et al., 2023; Škrinjar et al., 2025), undermining novelty, diversity, and data privacy. In this regard, another desirable property of generative models is their ability to generate high-quality, yet novel examples, known as generalisation.
@@
The matrix field $\widehat{\mathbf{\Gamma}}$ controls the local Dirichlet (carré du champ) energy, and can be efficiently and robustly estimated from data using diffusion geometry (jones2024diffusion; jones2024manifold), providing explicit geometric noise regularisation aligned with the data manifold. We demonstrate that across diverse synthetic and real-world datasets and neural network architectures, CDC-FM achieves comparable or better quality than FM, preserving the fine-grained details of the data, while substantially reducing memorisation and increasing generalisation. Our work provides a theoretical framework for the geometry-aware regularisation of flow-based generative models and a practical method that can be readily used in existing FM pipelines.
@@ 2 Background
In this paper, we are interested in modelling a set of data points $\mathbb{R}^{d}$ whose density $\nu(x)$ concentrates around an unknown lower-dimensional manifold. We begin by revisiting the standard FM formulation (lipman2023flow). We highlight that the FM probability path $p_{t}(x)$ induces at $t=1$ a homogeneous, isotropic Gaussian kernel approximation of $\nu$. This will motivate our generalised framework in Section 3, where we incorporate an anisotropic and inhomogeneous kernel approximation to strike a balance between faithfully modelling the data, substantially reducing memorisation, and improving generalisation. To illustrate the distinction between the two frameworks, we use a toy problem of learning to generate a circle from equidistant training points (Fig. 1).
@@ 2.1 Flow Matching
The velocity $u_{t}$, in turn, induces $p_{t}(x)$, the probability of finding a particle at $x$ at time $t$, via (2). Specifically, FM takes the final particle position $X_{1}\sim\nu$ as an auxiliary variable (a training point) and specifies the conditional flow $\psi(X|X_{1})$. Although the choice of $\psi(X|X_{1})$ has implications for the regularity of the density $\nu(x)$ (albergo_stochastic_2023), the standard choice is the affine flow
@@
$\mathcal{L}(\theta)=\mathbb{E}_{t,X_{1},X}||\hat{u}^{\theta}_{t}(X)-u_{t}(X|X_{1})||^{2}=\mathbb{E}_{t,X_{0},X_{1}}||\hat{u}^{\theta}_{t}(\psi(X_{0}|X_{1}))-\frac{d}{dt}\psi_{t}(X_{0}|X_{1})||^{2},$ (6)
@@ 3.1 Diffusive regularisation: Carré du champ Flow Matching (CDC-FM)
the displacement (optimal transport) interpolant between $\mu=\mathcal{N}(0,\mathbf{I})$ and an anisotropic Gaussian centred at $x_{1}\sim\nu$ that is geometrically aligned with the data manifold. While above we presented the special case of a Gaussian initial density $\mu$, we also derive a general flow path valid for arbitrary $\mu,\nu$ (Appendix A, Proposition 1). Note that as an alternative to (8) one could consider the naïve data augmentation that replaces the training points by perturbed samples, $x^{(i)}\mapsto\mathcal{N}(x^{(i)},\,\widehat{\mathbf{\Gamma}}(x^{(i)}))$, and then uses FM flow paths (5). However, as we prove, this approach is strictly different from (8) and generates suboptimal paths (Appendix B, Theorem 1).
@@
Our choice of flow path (8) provides an inductive bias in the FM model to learn the manifold. To see this, we substitute our flow path (8) into the FM loss (6), noting that the target velocity is $\tfrac{d}{dt}\psi_{t}(X_{0}|X_{1}=x_{1})=x_{1}+\big[\widehat{\mathbf{\Gamma}}(x_{1})^{1/2}-\mathbf{I}\big]X_{0}$. The random part (second term) can be approximated to leading order as $\big[\widehat{\mathbf{\Gamma}}(x_{1})^{1/2}-\mathbf{I}\big]X_{0}\sim\mathcal{N}(0,\mathbf{I}-2\widehat{\mathbf{\Gamma}}(x_{1})^{1/2}+\widehat{\mathbf{\Gamma}}(x_{1}))\simeq\mathcal{N}(0,\mathbf{I}-\widehat{\mathbf{\Gamma}}(x_{1}))$. This means that if $\widehat{\mathbf{\Gamma}}(x_{1})$ approximates the projection map onto the local tangent space, the dominant contributions to the velocity are approximately perpendicular (Fig. 1e), minimising tangential flows (Fig. 1c), which are associated with memorisation achilli_losing_2024.
@@
To further understand the geometric regularisation by the flow path (8), we may, as before, marginalise (9) over the target distribution $\nu(x_{1})$ to obtain
@@
$\nu\simeq p_{1}(x)=\frac{1}{N}\sum_{i=1}^{N}\mathcal{N}\left(x\middle|x^{(i)};\,\widehat{\mathbf{\Gamma}}(x^{(i)})\right),$ (10)
@@
replacing the FM approximation of $\nu$ in (7) with an anisotropic Gaussian mixture. As $\widehat{\mathbf{\Gamma}}(x_{1})$ approximates the projection map onto the local tangent space, we numerically observe that CDC-FM faithfully learns the data manifold (Fig. 1f) in contrast to FM (Fig. 1c).
@@
The approximation in (10) is equivalent to a noise-induced regularisation of data. We prove (Appendix C, Proposition 2) that the flow path (8) is equivalent to adding a geometry-aware anisotropic noise term to the continuity equation (2) to obtain a drift-diffusion process (22). The amount of smoothing introduced by diffusion integrated over the manifold is measured by the Dirichlet energy (32). Meanwhile, the integrand of the Dirichlet energy in (32) measuring the local smoothness around a training point $x^{(i)}$ is precisely $\widehat{\mathbf{\Gamma}}(x^{(i)})$, known as the carré du champ, justifying our construction. Compare Algorithms 1 and 2 in Appendix D for a condensed summary.
@@ 3.2 Estimating the carré du champ
To compute $\widehat{\mathbf{\Gamma}}$ that optimally captures the local geometry, we follow jones2024diffusion; jones2024manifold and provide a local kernel density estimate using the diffusion maps Laplacian (COIFMAN20065; berry2016variable). We compute a variable-bandwidth Gaussian kernel up to the $k$th neighbour of node $i$, $w_{\epsilon}(x^{(i)},x^{(j)})=\exp\!\left(-\|x^{(i)}-x^{(j)}\|^{2}/(\epsilon_{i}\epsilon_{j})\right)$, where $\epsilon_{i},\epsilon_{j}$ is the distance to the $k_{bw}$th nearest neighbour of $x^{(i)},x^{(j)}$, respectively. We use this to obtain a local estimate of the transition probabilities of a Markov process generating the data:
@@
$\widehat{\mathbf{\Gamma}}(x^{(i)})$ $=\mathbb{E}_{X\sim P_{i}}\left[\left(X-m^{*}(x^{(i)})\right)\left(X-m^{*}(x^{(i)})^{T}\right)\right],$ (12)
@@ 4.2 Quality-generalisation tradeoff for spatially heterogeneous data
To reinforce our results on the two-circles data, we considered animal motion capture data of the fruit fly, Drosophila melanogaster (deangelis). Points in this data are 31-frame 2D pose sequences of the 6-legged fly (Fig. 4a, inset), yielding a 372-dimensional ($2\times 6\times 31$) state space. The data manifold is parametrised by the walking speed on the longitudinal axis and phase on the cyclic coordinate, permitting visualisation in a 3D UMAP embedding (Fig. 4a).
@@
Amongst the models in Fig. 4b,c across different epochs, which one should one choose? When plotted against epochs, we see that generalisation in FM monotonically deteriorates (Fig. 4d). This means that FM models require early-stopping strategies to optimise. By contrast, CDC-FM test-set performance plateaus, meaning that early-stopping strategies are less relevant. Our results show that for a given sample quality, especially when high, there is a CDC-FM model with comparable or better generalisation and memorisation than FM, and vice versa.
The behaviours vary from highly stereotyped, tripod gaits to complex non-canonical gaits as the walking speed decreases (Fig. 4a). We expected that this heterogeneity would lead to spatially different trade-offs in quality, generalisation and memorisation. To visualise memorisation patterns, we took a late epoch (100k) for both FM and CDC-FM (Fig. 4b,c, black dots), finding that memorisation predominantly occurred on the right end of the UMAP (Fig. 4e). Memorisation correlated with the sparsity of the training points (Fig. 4f,g), estimated based on the distance to the closest neighbour, and occurred first at the sparsest points (Fig. 4g), corroborating our earlier findings (Fig. 3). By contrast, CDC-FM showed substantially lower (Fig. 4f,g) and less sparsity-dependent memorisation (Fig. 4g). Overall, these results indicate that geometry-aware regularisation is especially valuable when data varies heterogeneously over the data manifold. In the next section, we examine its limits as the intrinsic dimension and dataset size increase.
Figure 4: Quality-generalisation tradeoff for animal motion capture data. a Inset: an example fruit fly pose sequence. Point cloud, each point representing a 31-frame pose sequence, visualised in 3D UMAP coordinates. Shading indicates walking speed. b Generalisation against quality for CDC-FM for different $\widehat{\mathbf{\Gamma}}$ ranks, $d_{cdc}$, and for FM for different $\sigma_{\min}$. Black circles indicate epochs analysed in e. c Same as b, but for memorisation. d Generalisation against epochs. e Percentage of memorised samples nearest to a training point (FM: $\sigma_{\min}\!=\!0$, CDC-FM: $d_{cdc}=16$). f Variation of train data sparsity. g Average memorisation against sparsity for different epochs.

### v2

@@ Abstract
Deep generative models often face a fundamental tradeoff: high sample quality can come at the cost of memorisation, where the model reproduces training data rather than generalising across the underlying data geometry. We introduce Carré du champ flow matching (CDC-FM), a generalisation of flow matching (FM), that improves the quality-generalisation tradeoff by regularising the probability path with a geometry-aware noise. Our method replaces the homogeneous, isotropic noise in FM with a spatially varying, anisotropic Gaussian noise whose covariance captures the local geometry of the latent data manifold. We prove that this geometric noise can be optimally estimated from the data and is scalable to large data. Further, we provide an extensive experimental evaluation on diverse datasets (synthetic manifolds, point clouds, single-cell genomics, animal motion capture, and images) as well as various neural network architectures (MLPs, CNNs, and transformers). We demonstrate that CDC-FM consistently offers a better quality-generalisation tradeoff even when used as a latent space generation model. We observe significant improvements over standard FM in data-scarce regimes and in highly non-uniformly sampled datasets, which are often encountered in AI for science applications. Our work provides a mathematical framework for studying the interplay between data geometry, generalisation and memorisation in generative models, as well as a robust and scalable algorithm that can be readily integrated into existing flow matching pipelines.
@@ 1 Introduction
Deep generative models aim to sample from an unknown probability density $\nu(x)$, given finitely many training points $\mathcal{D}=\{x^{(i)}\}^{N}_{i=1}\subset\mathbb{R}^{d}$. Prominent paradigms include variational autoencoders (Kingma and Welling, 2014), generative adversarial networks (Goodfellow et al., 2014), diffusion processes (Sohl-Dickstein et al., 2015; Ho et al., 2020), and methods based on continuous normalising flows (Chen et al., 2019; Song and Ermon, 2019; Lipman et al., 2023; Albergo et al., 2023) (CNFs). Among these, CNFs have had striking recent success across domains from image generation to molecule design and weather prediction, owing to their ability to generate high-quality samples. However, a trivial and undesirable way of achieving high quality is through reproducing training points or close variants, known as memorisation. Thus, in addition to high quality, another desirable property of generative models is their ability to generate novel examples, known as generalisation.
Recently, there has been an increasing concern that the observed generative quality has been partly achieved through memorisation (Somepalli et al., 2023; Škrinjar et al., 2025), but possibly on large scales (Graber et al., 2025), undermining novelty, diversity, and data privacy. Recent works have shown that, geometrically, memorisation coincides with the sudden drop or complete vanishing of intrinsic dimensionality of the data manifold (Achilli et al., 2024; Ross et al., 2025; Ventura et al., 2025). In other words, during memorisation, the learned distribution degenerates towards an empirical measure supported on isolated training points, rather than a smooth, finite-dimensional manifold. This observation suggests that a way to address memorisation is by stabilising the intrinsic dimensionality and preserving non-degenerate tangent spaces.
@@
The matrix field $\widehat{\mathbf{\Gamma}}$ controls the local Dirichlet (carré du champ) energy, and can be efficiently and robustly estimated from data using diffusion geometry (jones2024diffusion; jones2024manifold; jones2026computing), providing explicit geometric noise regularisation aligned with the data manifold. We demonstrate that across diverse synthetic and real-world datasets and neural network architectures, CDC-FM achieves comparable or better quality than FM, preserving the fine-grained details of the data, while substantially reducing memorisation and increasing generalisation. Our work provides a theoretical framework for the geometry-aware regularisation of flow-based generative models and a practical method that can be readily used in existing FM pipelines.
@@ 2 Background
In this paper, we are interested in modelling a set of data points in $\mathbb{R}^{d}$ whose density $\nu(x)$ concentrates around an unknown lower-dimensional manifold. We begin by revisiting the standard FM formulation (lipman2023flow). We highlight that the FM probability path $p_{t}(x)$ induces at $t=1$ a homogeneous, isotropic Gaussian kernel approximation of $\nu$. This will motivate our generalised framework in Section 3, where we incorporate an anisotropic and inhomogeneous kernel approximation to strike a balance between faithfully modelling the data, substantially reducing memorisation, and improving generalisation. To illustrate the distinction between the two frameworks, we use a toy problem of learning to generate a circle from equidistant training points (Fig. 1).
@@ 2.1 Flow Matching
The velocity $u_{t}$, in turn, induces $p_{t}(x)$, the probability of finding a particle at $x$ at time $t$, via (2). Specifically, FM takes the final particle position $X_{1}\sim\nu$ as an auxiliary variable (a training point) and specifies the conditional flow $\psi_{t}(X|X_{1})$. Although the choice of $\psi_{t}(X|X_{1})$ has implications for the regularity of the density $\nu(x)$ (albergo_stochastic_2023), the standard choice is the affine flow
@@
$\mathcal{L}(\theta)=\mathbb{E}_{t,X_{1},X}||\hat{u}^{\theta}_{t}(X)-u_{t}(X|X_{1})||^{2}=\mathbb{E}_{t,X_{0},X_{1}}||\hat{u}^{\theta}_{t}(\psi_{t}(X_{0}|X_{1}))-\frac{d}{dt}\psi_{t}(X_{0}|X_{1})||^{2},$ (6)
@@ 3.1 Diffusive regularisation: Carré du champ Flow Matching (CDC-FM)
the displacement (optimal transport) interpolant between $\mu=\mathcal{N}(0,\mathbf{I})$ and an anisotropic Gaussian centred at $x_{1}\sim\nu$ that is geometrically aligned with the data manifold. In Appendix A (Proposition 1), we also derive the conditional probability paths valid for general $\mu,\nu$, illustrated in Fig. 1d.
@@
Note that as an alternative to (8) one could consider the naïve data augmentation that replaces the training points by perturbed samples, $x^{(i)}\mapsto\mathcal{N}(x^{(i)},\,\widehat{\mathbf{\Gamma}}(x^{(i)}))$, and then uses FM flow paths (5). However, as we prove, this approach generates strictly suboptimal paths (Appendix B, Theorem 1).
@@
Armed with the new conditional path (9), we may directly use the FM loss (6) to obtain the regularised velocity field. Compare Algorithms 1 and 2 in Appendix C for a condensed summary. A simple approximation shows that (8) provides an inductive bias for the learnt velocity field to approximate the data manifold. To see this, we substitute our flow path (8) into the FM loss (6), noting that the target velocity is $\tfrac{d}{dt}\psi_{t}(X_{0}|X_{1}=x_{1})=x_{1}+\big[\widehat{\mathbf{\Gamma}}(x_{1})^{1/2}-\mathbf{I}\big]X_{0}$. The random part (second term) can be approximated to leading order as $\big[\widehat{\mathbf{\Gamma}}(x_{1})^{1/2}-\mathbf{I}\big]X_{0}\sim\mathcal{N}(0,\mathbf{I}-2\widehat{\mathbf{\Gamma}}(x_{1})^{1/2}+\widehat{\mathbf{\Gamma}}(x_{1}))\simeq\mathcal{N}(0,\mathbf{I}-\widehat{\mathbf{\Gamma}}(x_{1}))$. This means that if $\widehat{\mathbf{\Gamma}}(x_{1})$ approximates the projection map onto the local tangent space, the dominant contributions to the velocity are approximately perpendicular (Fig. 1e), minimising tangential flows (Fig. 1c), which are associated with memorisation (achilli_losing_2024).
@@
To understand the mechanism behind the geometric regularisation induced by the flow path (8), we may first focus on the target distribution $\nu(x_{1})$. As before, we marginalise (9) to obtain
@@
$\nu\simeq p_{1}(x)=\frac{1}{N}\sum_{i=1}^{N}\mathcal{N}\left(x\middle|x^{(i)};\,\widehat{\mathbf{\Gamma}}(x^{(i)})\right).$ (10)
@@
We see that our flow path replaced the FM approximation of the data manifold $\nu$ in (7) with an anisotropic Gaussian mixture. As $\widehat{\mathbf{\Gamma}}(x_{1})$ approximates the projection map onto the local tangent space, we numerically observe that CDC-FM faithfully learns the data manifold (Fig. 1f) in contrast to FM (Fig. 1c).
@@
To obtain a deeper understanding of the regularisation mechanism, one may study marginal probability paths $p_{t}$. We prove (see Appendix D, Proposition 2) that our flow path (8) induces a geometry-aware anisotropic diffusion term to the continuity equation (2) to obtain a drift-diffusion process (22). For diffusion processes, one may use the Dirichlet energy (32) to quantify the smoothing introduced by the diffusion term arising from our flow paths. However, the Dirichlet energy is just the CDC field $\widehat{\mathbf{\Gamma}}(x^{(i)})$, measuring the local smoothness around a training point $x^{(i)}$, integrated over the data manifold. This justifies the construction of our flow path.
@@ 3.2 Estimating the carré du champ
To compute $\widehat{\mathbf{\Gamma}}$ that optimally captures the local geometry, we follow jones2024diffusion; jones2024manifold; jones2026computing and provide a local kernel density estimate using the diffusion maps Laplacian (COIFMAN20065; berry2016variable). We compute a variable-bandwidth Gaussian kernel up to the $k$th neighbour of node $i$, $w_{\epsilon}(x^{(i)},x^{(j)})=\exp\!\left(-\|x^{(i)}-x^{(j)}\|^{2}/(\epsilon_{i}\epsilon_{j})\right)$, where $\epsilon_{i},\epsilon_{j}$ is the distance to the $k_{bw}$th nearest neighbour of $x^{(i)},x^{(j)}$, respectively. We use this to obtain a local estimate of the transition probabilities of a Markov process generating the data:
@@
$\widehat{\mathbf{\Gamma}}(x^{(i)})$ $=\mathbb{E}_{X\sim P_{i}}\left[\left(X-m^{*}(x^{(i)})\right)\left(X-m^{*}(x^{(i)})\right)^{T}\right],$ (12)
@@ 4.2 Quality-generalisation tradeoff for spatially heterogeneous data
To reinforce our results on the two-circles data, we considered animal motion capture data of the fruit fly, Drosophila melanogaster (deangelis). Points in this data are 31-frame pose sequences where each snapshot represents the 2D position of the six legs relative to the body centroid (Fig. 4a, inset), yielding a 372-dimensional ($2\times 6\times 31$) state space. A 3D UMAP embedding of this data permits the visualisation of a data manifold, which is parametrised by the walking speed on the longitudinal axis and phase on the cyclic coordinate (Fig. 4a). We also performed a sensitivity analysis for the parameters $k_{bw}$ (bandwidth) and $k$ (number of nearest neighbours), finding that only $k$ had measurable influence on the estimation of the CDC field (Fig. A5). We obtained better generalisation and lower memorisation than FM for orders of magnitude changes of $k$. Nevertheless, optimal regularisation can be achieved for $k$ values large enough to provide an accurate kernel density estimate, but small enough to avoid short-cut connections in the manifold (Fig. A5).
Figure 4: Quality-generalisation tradeoff for animal motion capture data. a Inset: an example fruit fly pose sequence. Point cloud, each point representing a 31-frame pose sequence, visualised in 3D UMAP coordinates. Shading indicates walking speed. b Generalisation against quality for CDC-FM for different $\widehat{\mathbf{\Gamma}}$ ranks, $d_{cdc}$, and for FM for different $\sigma_{\min}$. Points are averaged over eight seeds. Black circles indicate epochs analysed in e. c Same as b, but for memorisation. d Generalisation and memorisation against epochs. Shaded error bars show one std from the mean. e Percentage of memorised samples nearest to a training point (FM: $\sigma_{\min}\!=\!0$, CDC-FM: $d_{cdc}=16$). f Variation of train data sparsity. g Average memorisation against sparsity for different epochs.
@@
Amongst the models in Fig. 4b,c across different epochs, which one should one choose? When plotted against epochs, we see that generalisation in FM monotonically deteriorates (Fig. 4d) and memorisation monotonically increases. This means that FM models require early-stopping strategies to optimise. By contrast, CDC-FM test-set performance plateaus and memorisation remains lower, meaning that early-stopping strategies are less relevant. Our results show that for a given sample quality, especially when high, there is a CDC-FM model with comparable or better generalisation and memorisation than FM, and vice versa.
The data manifold is not covered uniformly with data points. Rather, fast behaviours are generally more stereotypical and most densely represented (left side of manifold) than slow, complex movements (right side of manifold, Fig. 4a). We expected that this sparsity pattern would lead to varying trade-offs in quality, generalisation and memorisation. In visualising these patterns, we took a late epoch (100k) for both FM and CDC-FM (Fig. 4b,c, black dots), finding that memorisation predominantly occurred in sparse regions (Fig. 4e). Memorisation correlated with the sparsity of the training points (Fig. 4f,g), estimated based on the distance to the closest neighbour, and occurred first at the sparsest points (Fig. 4g), corroborating our earlier findings (Fig. 3). By contrast, CDC-FM showed substantially lower (Fig. 4f,g) and less sparsity-dependent memorisation (Fig. 4g). Overall, these results indicate that geometry-aware regularisation is especially valuable when data varies heterogeneously over the data manifold. In the next section, we examine its limits as the intrinsic dimension and dataset size increase.